# LLMShot: Reducing snapshot testing maintenance via LLMs

### v1

@@ Abstract
Snapshot testing has emerged as a critical technique for UI validation in modern software development, yet it suffers from substantial maintenance overhead due to frequent UI changes causing test failures that require manual inspection to distinguish between genuine regressions and intentional design changes. This manual triage process becomes increasingly burdensome as applications evolve, creating a need for automated analysis solutions. This paper introduces LLMShot, a novel framework that leverages vision-based Large Language Models to automatically analyze snapshot test failures through hierarchical classification of UI changes. To evaluate LLMShot’s effectiveness, we developed a comprehensive dataset using a feature-rich iOS application with configurable feature flags, creating realistic scenarios that produce authentic snapshot differences representative of real development workflows. Our evaluation using Gemma3 models demonstrates strong classification performance, with the 12B variant achieving over 84% recall in identifying failure root causes while the 4B model offers practical deployment advantages with acceptable performance for continuous integration environments. However, our exploration of selective ignore mechanisms revealed significant limitations in current prompting-based approaches for controllable visual reasoning. LLMShot represents the first automated approach to semantic snapshot test analysis, offering developers structured insights that can substantially reduce manual triage effort and advance toward more intelligent UI testing paradigms.
@@
snapshot testing, software testing, large language models, user interface testing, automated test analysis
@@ I Introduction
To address these challenges, we introduce LLMShot, a framework that leverages vision-based Large Language Models (LLMs) to automate snapshot test failure analysis. LLMShot employs Gemma3 models to categorize, quantify, and explain visual discrepancies using a hierarchical classification scheme. Unlike existing pass/fail approaches, our framework provides semantic understanding of UI changes, distinguishing between genuine regressions and intentional design modifications.
Our evaluation on 17 test cases demonstrates that LLMShot achieves strong classification performance, with the Gemma3:4b model reaching 76.47% hit rate and 66.67% F1-score. However, our exploration of ignore mechanisms revealed significant limitations in current prompting-based approaches.
@@
  • LLMShot, an intelligent snapshot analysis framework that performs semantic analysis and systematic categorization of visual discrepancies using a hierarchical classification scheme. This represents the first automated approach to snapshot test failure analysis reported in the academic literature, with a pipeline designed for CI/CD workflow integration.
  • A comprehensive dataset [^1] https://zenodo.org/records/15876540 with hierarchical ground truth labels for UI failure categorization, created using a feature-rich iOS demo app that demonstrates real-world features posing challenges for snapshot tests.
@@ II-A Visual GUI and Snapshot Testing
Early automated GUI testing focused on maximizing coverage through event-sequence generation and model-based exploration. Tools like Sapienz [5] used evolutionary algorithms for code coverage and crash discovery, while Stoat [6] built window-transition graphs to reduce redundant paths and combinatorial reducers [13] minimized event combinations for efficiency. ComboDroid [7] combined real-world traces with automated exploration for cost-effective fault detection. Video2Scenario (V2S) [14] automated test script generation through screen recording analysis using neural object detection, reproducing 89% of user actions. CAPdroid [15] enhanced V2S with improved detection and input validation while generating step descriptions from snapshots. More recently, WebExplor [16] and QExplore [17] formulated navigation as reinforcement learning tasks, improving lines executed, screens reached, and crashes detected beyond simple crawling. However, one notable gap in this body of work is the detection of visual regression that does not affect functional behavior but may still significantly impact user experience.
Lately, snapshot testing has gained popularity for validating UI consistency by comparing rendered states with baselines. However, it remains largely understudied in the literature. Bui and Rocha [18] compiled a JEST dataset from GitHub projects, capturing static attributes like stars, issues, and test files, but dynamic test execution remains unexplored. Cruz et al. [3] performed a grey literature review highlighting snapshot testing’s ease of use and regression-prevention benefits, while also noting its fragility and tendency toward false positives.
To address this, research explored smarter visual oracles. DiffDroid [10] executes Android apps across 130+ device profiles to flag inconsistencies with high precision, while dVermin [19] uses metamorphic approaches for layout issues under accessibility settings. Tools like Owl Eyes [11], Nighthawk [20], and GLIB [12] employ deep learning classifiers for visual glitches, achieving impressive precision (0.85 for mobile apps, 0.99 for game sprites) through synthetic data augmentation. However, while these tools can detect a wide range of visual defects, they fail to contextualize the changes, leaving developers without explanations for why the discrepancy occurred or whether it was intentional, which is essential in snapshot testing.
@@ II-B LLM-based Semantic Analysis for UI Verification
LLMs have introduced unprecedented potential for semantic analysis in software engineering, particularly for bug detection and UI verification. GPTDroid [8] frames testing as dialogue, achieving 32% improved activity coverage and 31% more bugs in 93 Android apps while helping developers fix 35 issues. VisionDroid [21] extends this to multimodal tasks, combining OCR with vision-language models to detect complex functional bugs like discrepancies between totals and itemized amounts.
While these advances represent significant progress, most LLM-based testing tools focus on bug detection and test generation, leaving snapshot testing largely unexplored. Traditional snapshot tests suffer from false positives or intentional UI modifications flagged as failures, accounting for up to 8.2% of commits [2] and requiring significant manual review effort. Existing guidelines recommend writing snapshot tests with "detailed and clear descriptions" and treating snapshots with the same care and discipline as code [3]. However, current tools provide no support for generating these detailed descriptions automatically for snapshot failures. LLMShot aims to alleviate this burden by using LLMs to semanticize the snapshot validation process, making it easier to distinguish between expected changes and true regressions.
III Dataset Creation
A key goal in snapshot testing is developing semantic understanding to distinguish intentional feature updates from genuine regressions. This requires realistic scenarios with meaningful explanatory labels beyond binary pass/fail results. While Bui et al. [18] developed what is, to the best of our knowledge, the first and only snapshot-specific dataset to date derived from 686 GitHub repositories, their work focuses on static analysis without executing tests or analyzing actual snapshot content, lacking infrastructure for test execution.
To address these limitations, we built our dataset using a real application with feature flags corresponding to realistic development scenarios: user privilege changes, content updates, and styling modifications. This generates authentic snapshot differences by toggling actual application features developers commonly modify—scenarios that fail snapshot tests but aren’t genuine regressions. Our dataset provides ground truth labels for change categories (Table I) classifying UI modifications from feature flag activation: visual changes (color, padding), content updates (text, data), structural modifications (layout), and behavioral changes (animations, semantic state).
TABLE I: Snapshot Test Failure Categories and Descriptions
Category Description
COLOR_CHANGE Different color values due to styling updates.
PADDING_CHANGE Change in margins/spacing; layout shifts.
UNKNOWN_<T> Change not from above categories; T names new reason.
For each record, we run twice, once with default settings and once with feature flags enabled, capturing snapshots at key UI states. This produces baseline and modified images per record. Each record contains: reference.png (baseline snapshot), failure.png (modified snapshot), diff.png (visual diff), and metadata.json (comprehensive context).
IV Methodology
The primary objective is distinguishing genuine UI regressions from intentional design changes and providing explanatory analysis for failing snapshot tests. Our approach employs LLMs to semantically analyze UI differences and categorize change types.
IV-A Core LLMShot
LLMShot processes snapshot test failures by leveraging vision-enabled LLMs. The core prompt is shown in Listing 1. Given a tuple of reference image, failure image, pixel-level diff, and predefined possible categories from Table I, LLMShot produces structured analysis outputs including change classification, quantitative measurements, affected UI elements, and natural language explanations.
@@ 3. DIFF: {difference} --- pixel-level diff image
Listing 1: Core LLM prompt framework of LLMShot
Change classification assigns label(s) from the possible categories or, if the LLM identifies changes that do not fit predefined categories, it can create new categories by giving a descriptive name for the newly identified change type.
The quantitative measurements include a pixel-wise difference score in a 0-1 scale, where 0 indicates identical images and 1 represents completely different images, and a semantic difference score in a 0-1 scale, where 0 means all expected components are present and correct, and 1 indicates all components are missing or incorrect. The affected UI elements output provides a list of specific component names that exhibit changes, while the natural language explanations offer an analysis of the key differences observed between the reference and failure snapshots.
IV-B Extension: Ignore Category Mechanism
In real-world scenarios, developers often need to ignore certain expected failure categories while focusing on others that indicate actual regressions. Developers can either ignore accepted failures identified by core LLMShot or directly specify ignore categories upfront, with the latter approach potentially enhancing the LLM’s sensitivity to detecting other categories. To address this need, we develop an ignore mechanism as an extension to LLMShot that allows developers to specify acceptable change categories, enabling focused analysis on unexpected differences that may indicate genuine issues.
The implementation modifies the core prompt (Listing 1) to instruct the LLM to disregard specified categories while maintaining sensitivity to other modifications, as shown in Listing 2.
@@ IV-B Extension: Ignore Category Mechanism
V Setup
V-A Dataset Construction
Our implementation uses a SwiftUI-based iOS application with tab-based architecture across Home, Profile, and Settings views, employing MVVM patterns with centralized state management. The app includes UI elements that change based on user state and active feature flags: content visibility, layout variations, and dynamic animations. We use runtime-configurable feature flags where each activates specific UI variations similar to code modifications. The combination of user states (free vs. premium) and feature flags creates diverse scenarios—the same flag produces different visual outcomes depending on user interface context, enabling comprehensive analysis with semantic understanding of UI changes.
We achieve snapshot capture by running tests using XCTest and XCUITest, extending default behaviors to support scrolling, element interactions, and app launching with specific configurations. Tests were executed using Xcode 16.3 on iPhone 15 Pro iOS 17.2. We use the swift-snapshot-testing library [22] to capture screenshots, running each test twice as described in the methodology.
@@ V-A Dataset Construction
We run a total of 17 tests with 19 different flags enabled. Two tests are multi-flag, having two flags enabled simultaneously. The distribution of flag categories across the 17 tests shows that COLOR_CHANGE is the most prevalent type of UI discrepancy, occurring in 5 tests. Three categories—PADDING_CHANGE, CONTENT_CHANGE, and LAYOUT_CHANGE—each appear in 3 tests. TEXT_CHANGE and ANIMATION_PHASE each occur in 2 tests, while ANIMATION_CHANGE is the least frequent, appearing in only 1 test.
@@
The dataset exhibits a mean pixel difference of 0.053 ± 0.054 (mean ± std) across all test cases, calculated using Equation 1. This indicates that, on average, the reference and failure images differ by approximately 5.3% at the pixel level. The substantial variation across test cases demonstrates dataset diversity. We quantify image-level differences using a normalized pixel-wise difference score:
@@
V-B LLMShot
We evaluate LLMShot using Gemma3 [23] models in 4B and 12B parameter variants with a temperature of 0.1 through the Ollama framework [24]. The core prompt framework (Listing 1) guides the LLM to analyze reference, failure, and diff images systematically, producing structured JSON responses.
V-C Evaluation Metrics
@@
VI Evaluation
VI-A RQ1: How does LLMShot perform in Failure Classification?
The primary objective is to establish whether visual language models can accurately categorize different types of UI changes in snapshot test failures, providing developers with structured explanations of detected differences. To evaluate LLMs’ automated snapshot failure analysis capability, we applied the core LLMShot prompt framework to our dataset.
Table II shows LLMShot’s classification performance across model sizes. The larger gemma3:12b model outperformed the smaller gemma3:4b model across all metrics, achieving over 80% Hit Rate and Recall. However, Precision was notably lower for both models, indicating some incorrect predictions. F1-Scores of 74.42% and 66.67% respectively demonstrate balanced performance when accounting for false positives. Both models showed conservative prediction behavior, averaging fewer than 1.5 labels per test.
Finding 1: Our results indicate that LLMShot demonstrates strong potential in automatically classifying snapshot test failure root causes on our dataset.
TABLE II: Performance of Core LLMShot
@@ VI-A RQ1: How does LLMShot perform in Failure Classification?
  • Hit Rate: Test Case Hit Rate
  • Semantic Diff.: Semantic Difference
VI-B RQ2: How does the ignore-category mechanism affect performance of LLMShot?
In real-world development scenarios, developers must selectively ignore failure categories corresponding to intentional changes while maintaining detection of categories that indicate actual regressions. One approach is to use core LLMShot, ignore the intended classes, and highlight other classes from its results, since LLMs can provide multiple labels. The other approach is to use an extension of the ignore category mechanism. Here, we investigate whether LLMShot can selectively ignore specified UI change categories while maintaining sensitivity to other change types.
To evaluate this extension on our dataset, we follow two strategies: "Ignore From Analysis" runs the core LLMShot once and then runs the extension by instructing the LLM to ignore the first category identified by core LLMShot, and "Ignore From Ground Truth" runs the extension while ignoring the first ground truth category. In both approaches, if LLMShot outputs the ignored category, it is considered a false positive. When the ignored category represents the only correct classification, the expected LLMShot response is to return an empty set indicating no other categories found, and this is considered a hit.
For this experiment, we used the smaller Gemma3:4b model due to its suitability for CI/CD deployment and practical implementation constraints. Table III reveals significant performance degradation across all ignore strategies compared to core LLMShot. "Ignore From Analysis" performed worst, with Hit Rate dropping to 18.75%, Recall to 28.57%, and F1-Score plummeting to 13.79%. "Ignore From Ground Truth" showed better recall but maintained the same low precision (9.09%), resulting in only marginal F1-Score improvement to 16.67%. Both ignore strategies demonstrated poor compliance rates (31.25% and 35.29% respectively), indicating LLMShot frequently predicted ignored categories despite explicit instructions. Additionally, pixel difference errors increased substantially (0.157±0.244 and 0.287±0.339 vs 0.095±0.161 baseline), and unknown classification rates rose, particularly for "Ignore From Analysis" (25% vs 11.76% baseline).
Finding 2: Current LLM-based ignore mechanisms struggle severely with selective change filtering, showing substantial performance degradation across all classification metrics and poor instruction compliance compared to core LLMShot analysis.

### v2

@@ Abstract
Snapshot testing has emerged as a critical technique for UI validation in modern software development, yet it suffers from substantial maintenance overhead due to frequent UI changes causing test failures that require manual inspection to distinguish between genuine regressions and intentional design changes. This manual triage process becomes increasingly burdensome as applications evolve, creating a need for automated analysis solutions. This paper introduces LLMShot, a novel framework that leverages Vision-Language Models (VLMs) to automatically analyze snapshot test failures through semantic classification of UI changes. To evaluate LLMShot’s effectiveness, we developed a comprehensive dataset using a feature-rich iOS application with configurable feature flags, creating realistic scenarios that produce authentic snapshot differences representative of real development workflows. Our evaluation using Gemma3 models demonstrates strong classification performance, with the 12B variant achieving over 84% recall in identifying failure root causes while the 4B model offers practical deployment advantages with acceptable performance for continuous integration environments. However, our exploration of selective ignore mechanisms revealed significant limitations in current prompting-based approaches for controllable visual reasoning. LLMShot represents the first automated approach to semantic snapshot test analysis, offering developers structured insights that can substantially reduce manual triage effort and advance toward more intelligent UI testing paradigms.
@@
snapshot testing, software testing, large language models, vision-language models, user interface testing, automated test analysis
@@ I Introduction
To address these challenges, we introduce LLMShot, a framework that leverages VLMs to automate snapshot test failure analysis. LLMShot employs Gemma3 models to categorize, quantify, and explain visual discrepancies through semantic classification into a predefined taxonomy. Unlike existing pass/fail approaches, our framework provides semantic understanding of UI changes, distinguishing between genuine regressions and intentional design modifications.
Our evaluation demonstrates strong classification performance, with the Gemma3 12B model achieving 84.21% recall. The more efficient 4B model also performs robustly, reaching 78.95% recall. However, our exploration of the ignore mechanism reveals that current prompting techniques are insufficient for complex controllable reasoning, highlighting a key challenge for future research.
@@
  • The design and implementation of LLMShot, the first automated framework for the semantic analysis of snapshot test failures. LLMShot categorizes visual discrepancies according to a structured taxonomy of change types and through its structured JSON output, provides a pipeline suitable for CI/CD integration.
  • A new, publicly available dataset [^1] https://zenodo.org/records/15876539 for snapshot testing research. Our dataset provides ground truth labels based on our UI failure taxonomy and contains realistic and challenging UI changes generated from a feature-rich iOS application.
@@ II-A Visual GUI and Snapshot Testing
Early automated GUI testing aimed to maximize functional coverage. This foundational work includes tools for event-sequence generation and exploration like Sapienz [5] and Stoat [6], efficiency improvements through techniques like combinatorial reduction [13], and automated test script generation from recordings or traces [7, 14, 15]. More recently, reinforcement learning has been applied to improve navigation and bug detection [16, 17]. While comprehensive, this research primarily targets functional correctness, leaving the distinct challenge of non-functional visual regressions largely unaddressed.
Snapshot testing has gained popularity precisely to fill this gap by comparing rendered UI states against baselines. However, it remains largely understudied in the literature. Existing studies are often preliminary; for instance, Bui and Rocha [18] compiled a dataset of static project attributes without executing tests, while a grey literature review by Cruz et al. [3] noted the technique’s benefits but also its fragility and high rate of false positives.
In parallel with the rise of snapshot testing, broader research has explored smarter visual oracles to address the general fragility of visual UI validation. These include tools for detecting rendering inconsistencies across devices like DiffDroid [10], layout issues using metamorphic testing [19], and visual glitches via deep learning classifiers [11, 20, 12]. However, a critical limitation unites these advanced oracles: while they excel at detecting visual defects, they do not explain them. They fail to provide the semantic context needed to determine if a failure is a genuine regression or an intentional change.
@@ II-B LLM-based Semantic Analysis for UI Verification
The advent of Vision-Language Models (VLMs) has unlocked new potential for semantic analysis in UI testing. Early successes like GPTDroid [8] and VisionDroid [21] have already leveraged VLMs to find complex functional bugs and improve test generation by reasoning about on-screen content.
Despite these advances, their application has been concentrated on functional bug detection, leaving the core problem of snapshot test analysis unexplored. The primary challenge in snapshot testing is not detection, but the high manual effort required to triage failures caused by intentional UI changes [2]. While guidelines recommend writing tests with “detailed and clear descriptions” [3], no tools provide this automatically. LLMShot is the first work to bridge this explanatory gap. It applies VLMs to semanticize the validation process itself, providing the contextual analysis needed to distinguish intentional changes from true regressions.
III Methodology
Our primary objective is to distinguish genuine UI regressions from intentional design changes by providing automated, explanatory analysis for failing snapshot tests. To this end, we introduce the LLMShot framework, which we evaluate on a new dataset designed specifically for this task.
III-A LLMShot
LLMShot processes snapshot test failures using VLMs. The core of the framework is a structured prompt, shown in Listing 1, which was developed through an iterative design process.
@@ 3. DIFF: {difference} --- pixel-level diff image
Listing 1: LLM prompt framework of LLMShot
Following the instructions in this prompt, LLMShot operates on a specific set of inputs to produce a structured analysis. The primary input is a tuple of the three images referenced in the prompt (before, after, and diff). The corresponding output is the single JSON object specified in the prompt.
The JSON output provides a multi-faceted analysis. The Change Classification assigns one or more labels from the predefined categories (Table I). A key feature is that if the model identifies a change that does not fit the existing taxonomy, it can dynamically create a new category by giving a descriptive name to the newly identified change type. The Quantitative Measurements include a pixel-wise difference score (on a 0-1 scale where 0 is identical) and a semantic difference score (on a 0-1 scale where 0 means all expected components are correct). Finally, the output provides a list of specific Affected UI Elements and a concise Natural Language Explanation analyzing the key differences observed between the snapshots.
III-B Extension: Ignore Category Mechanism
In practice, developers often need to ignore expected changes (e.g., a planned color update) to focus on unexpected ones that may be genuine regressions. To simulate this critical workflow, we developed an “ignore mechanism” extension. The implementation modifies the core prompt by appending an instruction for the VLM to disregard a specified category while maintaining its sensitivity to any other modifications. The template for this extension is shown in Listing 2.
@@ III-B Extension: Ignore Category Mechanism
III-C Evaluation Dataset
A key goal in snapshot testing is developing semantic understanding to distinguish intentional feature updates from genuine regressions. This requires realistic test scenarios with meaningful labels that go beyond a binary pass/fail result. The only prior snapshot-specific dataset, from Bui et al. [18] provides a valuable foundation of static project analysis, but lacks the executable tests and visual snapshot content essential for analyzing failure root causes.
To fill this critical gap, we introduce a new, publicly available dataset designed specifically for this task. We generate this dataset from a purpose-built iOS application, using its integrated feature flags to produce authentic UI changes that mimic common development scenarios. Each resulting failure instance is then manually classified by two authors and labeled according to our detailed taxonomy of UI changes (Table I). This taxonomy covers a wide range of modifications, including visual changes (e.g., color, padding), content updates (e.g., text, data), structural modifications (e.g., layout), and behavioral changes (e.g., animations).
This process results in a novel dataset containing not only baseline and failure snapshots but also rich contextual metadata and corresponding ground truth labels. To the best of our knowledge, it is the first publicly available dataset to provide these dynamic artifacts, thereby enabling new opportunities for the rigorous, empirical evaluation of tools for semantic snapshot test analysis.
TABLE I: Snapshot Test Failure Categories and Descriptions
Category Description
COLOR_CHANGE Different color values due to styling updates.
PADDING_CHANGE Change in margins/spacing; layout shifts.
UNKNOWN_<T> Change not from above categories; T names new reason.
IV Setup
To evaluate LLMShot, we detail the construction of our dataset, our model configuration, and the metrics used for assessment.
IV-A Dataset Construction
We generate our dataset from a purpose-built, feature-rich iOS application, which we develop in SwiftUI specifically for this study. We design the app with a non-trivial architecture (MVVM, tab-based navigation) and dynamic UI elements that change based on user state and active feature flags to provide a realistic evaluation context.
Data Generation. We use the application’s 19 runtime-configurable feature flags, a standard industry practice that allows for the controlled creation of diverse UI modifications such as changes to content visibility, layout, and dynamic animations. The combination of these flags with different user states (free vs. premium) creates diverse scenarios where the same flag can produce different visual outcomes depending on the UI context. We execute each of our 17 test cases using standard XCTest and XCUITest suites, with extensions to support scrolling and element interactions. For capturing the images themselves, we use the swift-snapshot-testing library [22], running each test twice: once with default settings to capture a baseline, and once with a feature flag enabled to produce a failure. Two tests involve multi-flag activations to create more complex scenarios. The test execution environment is Xcode 16.3 on an iPhone 15 Pro running iOS 17.2.
Ground Truth Labeling. To establish a reliable ground truth, two of the authors independently classify each resulting failure according to the categories in Table I. They then resolve all disagreements through a joint discussion to finalize the labels.
@@ IV-A Dataset Construction
Dataset Characteristics. The distribution of flag categories across the 17 tests shows that COLOR_CHANGE is the most prevalent type of UI discrepancy, occurring in 5 tests. Three categories—PADDING_CHANGE, CONTENT_CHANGE, and LAYOUT_CHANGE—each appear in 3 tests. TEXT_CHANGE and ANIMATION_PHASE each occur in 2 tests, while ANIMATION_CHANGE is the least frequent, appearing in only 1 test.
@@
The dataset exhibits a mean pixel difference of 0.053 $\pm$ 0.054 (mean $\pm$ std) across all test cases, calculated using Equation 1. This indicates that, on average, the reference and failure images differ by approximately 5.3% at the pixel level. The substantial variation across test cases demonstrates dataset diversity. We quantify image-level differences using a normalized pixel-wise difference score:
@@
IV-B Model Configuration
We select the Gemma3 [23] family of models for our evaluation for two key reasons. First, they are state-of-the-art, open-weights models that support multi-image prompts, a critical feature for our framework which requires analyzing reference, failure, and diff images simultaneously. Second, their availability in multiple sizes allows us to analyze performance-to-scale trade-offs. We evaluate LLMShot using the 4B and 12B parameter variants with a temperature of 0.1 through the Ollama framework [24]. The prompt framework (Listing 1) guides the LLM to analyze reference, failure, and diff images systematically, producing structured JSON responses.
IV-C Evaluation Metrics
@@
V Evaluation
V-A RQ1: How does LLMShot perform in Failure Classification?
The primary objective is to establish whether visual language models can accurately categorize different types of UI changes in snapshot test failures. To evaluate this capability, we applied the LLMShot prompt framework to our dataset using both Gemma3-4B and Gemma3-12B models.
Table II shows LLMShot’s classification performance across model sizes. The larger Gemma3-12B model outperformed the smaller Gemma3-4B model across all metrics, achieving over 80% Hit Rate and 84.21% Recall. However, Precision was notably lower for both models, indicating a tendency to generate some incorrect predictions alongside correct ones. F1-Scores of 74.42% and 66.67% respectively demonstrate balanced performance when accounting for these false positives. The superiority of the larger model is further evidenced by its lower ‘Unknown Rate‘ (5.88% vs 11.76%) and more accurate pixel difference estimation. Both models showed conservative prediction behavior, averaging fewer than 1.5 labels per test.
Finding 1: Our results indicate that LLMShot demonstrates strong potential in automatically classifying snapshot test failure root causes. The larger 12B model shows superior performance in both classification and quantitative reasoning, though both models exhibit a trade-off between high recall and moderate precision.
TABLE II: Performance of LLMShot
@@ V-A RQ1: How does LLMShot perform in Failure Classification?
  • Hit Rate: Test Case Hit Rate, Semantic Diff.: Semantic Difference
V-B RQ2: How does the ignore-category mechanism affect the performance of LLMShot?
In real-world development, a critical task is to selectively ignore failures from intentional changes while still detecting regressions. We investigate whether our prompt-based ignore mechanism can achieve this, using the more efficient Gemma3:4b model. We evaluate two strategies: “Ignore From Analysis” (IFA), which instructs the model to ignore the category it first identified, and “Ignore From Ground Truth” (IFGT), which instructs it to ignore a known ground truth category.
The results, presented in Table III, reveal a significant performance degradation across both strategies. The IFA strategy performed worst, with its Hit Rate dropping to 18.75% and its F1-Score decreasing sharply to 13.79%. The IFGT strategy presented a more complex outcome. While its Recall reached 100.00%, this result is misleading as it was coupled with an extremely low precision of 9.09%. This suggests the model adopted a noisy prediction strategy, which is further reflected by its marginal F1-Score of only 16.67%.
Furthermore, both strategies demonstrated poor instruction compliance, with Ignore Compliance Rates of only 31.25% for IFA and 35.29% for IFGT. This indicates that the model frequently failed to disregard the specified ignore category. This difficulty is also reflected in other metrics: pixel difference errors increased substantially (0.157$\pm$0.244 for IFA and 0.287$\pm$0.339 for IFGT vs 0.095$\pm$0.161 baseline), and the unknown classification rate for IFA rose to 25%.
Finding 2: Current prompt-based ignore mechanisms struggle severely with selective change filtering. They exhibit poor instruction compliance and suffer a substantial degradation across all classification metrics. The high recall observed in one strategy is an artifact of its extremely low precision, rendering the approach unreliable for practical use.